# Review of the low-lying excited baryons $\Sigma^{*}(1/2^{-})$

### v1

@@ I Introduction
As composite subatomic particles made of quarks and gluons held together by the strong interaction, hadrons are categorized into $q\bar{q}$ mesons and $qqq$ baryons within the conventional quark model proposed by Gell-Mann and Zweig Gell-Mann:1964ewy ; Zweig:1964ruk . So far, all the ground-state mesons and baryons composed of light quarks ($u,d,s$) have been well established, and a huge number of excited light mesons and baryons have been observed experimentally ParticleDataGroup:2022pth , some of which can not be well described in the conventional quark models. For instance, recently BESIII has observed one isoscalar state $\eta_{1}(1855)$ with exotic quantum numbers $J^{PC}=1^{-+}$ in the process $J/\psi\to\gamma\eta\eta^{\prime}$, which could be a candidate for hybrid hadrons BESIII:2022riz ; Chen:2022qpd . Those exotic light hadrons challenge our understanding of the quark model and provide insights into the strong interaction dynamics within Quantum Chromodynamics (QCD) Chen:2016spr ; Chen:2022asf ; Guo:2017jvc ; Oset:2016lyh ; Liu:2024uxn ; Gao:1999ar ; Hall:2014uca .
There are some puzzles for the low-lying excited baryons with the quantum numbers of spin-parity $J^{P}=1/2^{-}$. One puzzle is the “mass reverse problem”. In the naive quark model, the mass of the $N(1535)$ with spin-parity quantum numbers $J^{P}=1/2^{-}$ should be lower than the one of the radial excitation $N(1440)$ with $J^{P}=1/2^{+}$, and the one of the $\Lambda(1405)$ ($J^{P}=1/2^{-}$) with strangeness $S=-1$ ParticleDataGroup:2022pth ; Liu:2005pm ; Geng:2008cv ; Capstick:1986ter , as shown in Fig. 1. Meanwhile, the $N(1535)$ strongly couples to the channels with strangeness, such as the $\eta N$, $K\Lambda$, and $K\Sigma$, which implies that this state has significant sea-quark admixtures. The low-lying excited baryon $\Lambda(1670)$ exhibits itself as an enhancement in the $\eta\Lambda$ invariant mass distribution of the processes $K^{-}p\to\eta\Lambda$ CrystalBall:2001uhc and $\Lambda_{c}^{+}\to\eta\Lambda\pi^{+}$ BESIII:2018qyg ; Belle:2020xku , and as a dip structure around the $\eta\Lambda$ threshold in the $\bar{K}^{0}n$ invariant mass distribution of the process $K^{-}p\to\bar{K}^{0}n$ Gopal:1976gs . Recently, the Belle Collaboration observed a cusp structure around the $\eta\Lambda$ threshold in the $pK^{-}$ invariant mass distribution of the process $\Lambda_{c}^{+}\to pK^{-}\pi^{+}$, which could be associated with the $\Lambda(1670)$ Belle:2022cbs ; Zhang:2024jby . Thus, these exotic properties of the low-lying excited baryons with the quantum numbers of spin-parity $J^{P}=1/2^{-}$ are difficult to explain in the simple quenched quark model and should have a much richer structure than the three valence quark composition Zou:2010tc .
@@
In a word, the $\Sigma^{*}(1/2^{-})$ with a mass around 1.4 GeV as a partner of $\Lambda(1405)$ has not been found yet. Since searching for the state $\Sigma^{*}(1/2^{-})$ is crucial to deepening our understanding of low-lying excited baryons Crede:2013kia ; Klempt:2009pi ; Oset:2016lyh ; Chen:2022zgm ; Ramos:2002xh and there are many theoretical works have been devoted to this issue, we would like to review the experimental and theoretical progress on the studies of the low-lying excited baryon $\Sigma^{*}(1/2^{-})$.
@@ II.1 $\Sigma(1620)$
The $\Sigma(1620)$ with $J^{P}=1/2^{-}$ was listed as a 2-star resonance in the previous version of PRR and downgraded to 1-star in the present version ParticleDataGroup:2022pth . Although its mass $(1600\sim 1650)$ MeV is consistent with the predicted mass around 1650 NeV of the quenched quark models Capstick:1986ter , the evidence for its existence is very weak Zou:2016bxw . Based on the multi-channel analysis of the $\bar{K}N$ reactions, both Ref. Kim:1971zxa and Ref. Langbein:1972uhb claimed the evidence for a $\Sigma^{*}(1/2^{-})$ with a mass around 1620 MeV, but predicted different branching ratios for this resonance. However, Ref. Baillon:1975sk found no signal for $\Sigma^{*}(1/2^{-})$ between 1600 MeV and 1650 MeV through the analysis of the reaction $\bar{K}N\to\Lambda\pi$, and suggested the existence of the $\Sigma(1660)$ with $J^{P}=1/2^{-}$, supported by the multi-channel analysis of the $\bar{K}N$ reactions Kim:1971zxa ; Langbein:1972uhb . In Ref. Carroll:1976gs , the analysis of the total cross sections for $K^{-}p$ and $K^{-}n$ with all proper final states indicated some $\Sigma$ resonances near 1600 MeV with quantum numbers undetermined. In Ref. Morris:1978ia , the authors analyzed the reaction $K^{-}n\to\pi^{-}\Lambda$ and obtained two solutions, with one solution indicating a $\Sigma^{*}(1/2^{-})$ near 1600 MeV, and the other showing no resonant structure below the $\Sigma(1670)$ ($J^{P}=3/2^{-}$).
@@ III.6 Charmed decays $\Lambda_{c}^{+}\to\pi^{+}\pi^{0}\pi^{-}\Sigma^{+}$
Since triangle singularity plays an important role in many processes Liu:2015taa ; Bayar:2016ftu ; Liang:2019jtr ; Dai:2018hqb ; Wang:2016dtb ; Guo:2019twa , Ref. Xie:2018gbi studied the $\Sigma^{*}(1/2^{-})$ production in the decay of $\Lambda_{c}^{+}\to\pi^{+}\pi^{0}\pi^{-}\Sigma^{+}$ by considering the triangle loop, which develops a singularity of the invariant mass of $\pi^{0}\pi^{-}\Sigma^{+}$ system around 1880 MeV. Here, the $\Sigma^{*}(1/2^{-})$ with a mass around 1430 MeV is a dynamically generated resonance from the meson-baryon interaction and is a cusp structure Roca:2013cca . It is found that a narrow peak, tied to the $\Sigma^{*}(1/2^{-})$ state, appears in the final $\pi^{-}\Sigma^{+}$ mass spectrum at the energy around the $\bar{K}N$ threshold. Meanwhile, the branching ratio for the $\Sigma^{*}(1/2^{-})$ signal is predicted to be of the order of $10^{-4}$ Xie:2018gbi . In the future, Belle II, BESIII, and the future Super Tau-Charm Factory (STCF) experiments could test such predictions.
@@ Acknowledgements
We would like to thank Eulogio Oset for useful comments. LSG and JJX are partly supported by the National Key R&D Program of China under Grant No. 2023YFA1606700. This work is supported by the Natural Science Foundation of Henan under Grant No. 232300421140 and No. 222300420554, the National Natural Science Foundation of China under Grant No. 12192263, 12205075, 12175239, 12221005, 12075288, 12361141819, and also by the Project of Youth Backbone Teachers of Colleges and Universities of Henan Province (2020GGJS017), and also by the Open Project of Guangxi Key Laboratory of Nuclear Physics and Nuclear Technology, No. NLK2021-08, and also by the Central Government Guidance Funds for Local Scientific and Technological Development, China (No. Guike ZY22096024), and also by the National Key Research and Development Program of China under Contracts 2020YFA0406400, and also by the Chinese Academy of Sciences under Grant No. YSBR-101, and also by the Youth Innovation Promotion Association CAS.

### v2

@@ I Introduction
As composite subatomic particles made of quarks and gluons held together by the strong interaction, hadrons are categorized into $q\bar{q}$ mesons and $qqq$ baryons within the conventional quark model proposed by Gell-Mann and Zweig Gell-Mann:1964ewy ; Zweig:1964ruk . So far, all the ground-state mesons and baryons composed of light quarks ($u,d,s$) have been well established, and a huge number of excited light mesons and baryons have been observed experimentally ParticleDataGroup:2022pth , some of which can not be well described in the conventional quark models. For instance, recently BESIII has observed one isoscalar state $\eta_{1}(1855)$ with exotic quantum numbers $J^{PC}=1^{-+}$ in the process $J/\psi\to\gamma\eta\eta^{\prime}$, which could be a candidate for hybrid hadrons BESIII:2022riz ; Chen:2022qpd . Those exotic hadrons challenge our understanding of the quark model and provide insights into the strong interaction dynamics within Quantum Chromodynamics (QCD) Chen:2016spr ; Chen:2022asf ; Guo:2017jvc ; Oset:2016lyh ; Liu:2024uxn ; Gao:1999ar ; Hall:2014uca ; Zhang:2024ery ; Shi:2023ntq ; Wang:2023vtx ; Wang:2022xga ; Xie:2024wbd .
There are some puzzles for the low-lying excited baryons with the quantum numbers of spin-parity $J^{P}=1/2^{-}$. One puzzle is the “mass reverse problem”. In the naive quark model, the mass of the $N(1535)$ with spin-parity quantum numbers $J^{P}=1/2^{-}$ should be lower than the one of the radial excitation $N(1440)$ with $J^{P}=1/2^{+}$, and the one of the $\Lambda(1405)$ ($J^{P}=1/2^{-}$) with strangeness $S=-1$ ParticleDataGroup:2022pth ; Liu:2005pm ; Geng:2008cv ; Capstick:1986ter , as shown in Fig. 1. Although there are some debates on the relative ratio between the coupling constants of $N(1535)N\eta$ and $N(1535)K\Lambda$ Liu:2005pm ; Geng:2008cv ; Mart:2013ida ; Wang:2022osj obtained in different ways, they all indicate that the $N(1535)$ strongly couples to the channels with strangeness, which implies that this state has a significant penta-quark component with hidden strangeness. Concerning another low-lying $J^{P}=1/2^{-}$ excited baryon $\Lambda(1670)$, it exhibits itself as an enhancement in the $\eta\Lambda$ invariant mass distribution of the processes $K^{-}p\to\eta\Lambda$ CrystalBall:2001uhc and $\Lambda_{c}^{+}\to\eta\Lambda\pi^{+}$ BESIII:2018qyg ; Belle:2020xku , and as a dip structure around the $\eta\Lambda$ threshold in the $\bar{K}^{0}n$ invariant mass distribution of the process $K^{-}p\to\bar{K}^{0}n$ Gopal:1976gs . Recently, the Belle Collaboration observed a cusp structure around the $\eta\Lambda$ threshold in the $pK^{-}$ invariant mass distribution of the process $\Lambda_{c}^{+}\to pK^{-}\pi^{+}$, which could be associated with the $\Lambda(1670)$ Belle:2022cbs ; Zhang:2024jby . Thus, these exotic properties of the low-lying excited baryons with the quantum numbers of spin-parity $J^{P}=1/2^{-}$ are difficult to explain in the simple quenched quark model and should have a much richer structure than the three valence quark composition Zou:2010tc .
@@
In a word, the $\Sigma^{*}(1/2^{-})$ with a mass around 1.4 GeV as a partner of $\Lambda(1405)$ has not been found yet. Since searching for the state $\Sigma^{*}(1/2^{-})$ is crucial to deepening our understanding of low-lying excited baryons Crede:2013kia ; Klempt:2009pi ; Oset:2016lyh ; Chen:2022zgm ; Ramos:2002xh ; Azimov:2003bb ; Azimov:1970ei and there are many theoretical works have been devoted to this issue, we would like to review the experimental and theoretical progress on the studies of the low-lying excited baryon $\Sigma^{*}(1/2^{-})$.
@@ II.1 $\Sigma(1620)$
The $\Sigma(1620)$ with $J^{P}=1/2^{-}$ was listed as a 2-star resonance in the previous version of PRR and downgraded to 1-star in the present version ParticleDataGroup:2022pth . Although its mass $(1600\sim 1650)$ MeV is consistent with the predicted mass around 1650 MeV of the quenched quark models Capstick:1986ter , the evidence for its existence is very weak Zou:2016bxw . Based on the multi-channel analysis of the $\bar{K}N$ reactions, both Ref. Kim:1971zxa and Ref. Langbein:1972uhb claimed the evidence for a $\Sigma^{*}(1/2^{-})$ with a mass around 1620 MeV, but predicted different branching ratios for this resonance. However, Ref. Baillon:1975sk found no signal for $\Sigma^{*}(1/2^{-})$ between 1600 MeV and 1650 MeV through the analysis of the reaction $\bar{K}N\to\Lambda\pi$, and suggested the existence of the $\Sigma(1660)$ with $J^{P}=1/2^{-}$, supported by the multi-channel analysis of the $\bar{K}N$ reactions Kim:1971zxa ; Langbein:1972uhb . In Ref. Carroll:1976gs , the analysis of the total cross sections for $K^{-}p$ and $K^{-}n$ with all proper final states indicated some $\Sigma$ resonances near 1600 MeV with quantum numbers undetermined. In Ref. Morris:1978ia , the authors analyzed the reaction $K^{-}n\to\pi^{-}\Lambda$ and obtained two solutions, with one solution indicating a $\Sigma^{*}(1/2^{-})$ near 1600 MeV, and the other showing no resonant structure below the $\Sigma(1670)$ ($J^{P}=3/2^{-}$).
@@ III.6 Charmed decays $\Lambda_{c}^{+}\to\pi^{+}\pi^{0}\pi^{-}\Sigma^{+}$
Since triangle singularity plays an important role in many processes Liu:2015taa ; Bayar:2016ftu ; Liang:2019jtr ; Dai:2018hqb ; Wang:2016dtb ; Guo:2019twa , Ref. Xie:2018gbi studied the $\Sigma^{*}(1/2^{-})$ production in the decay of $\Lambda_{c}^{+}\to\pi^{+}\pi^{0}\pi^{-}\Sigma^{+}$ by considering the triangle loop, which develops a singularity of the invariant mass of $\pi^{0}\pi^{-}\Sigma^{+}$ system around 1880 MeV. It is found that a cusp structure, tied to the $\Sigma^{*}(1/2^{-})$ state, appears in the final $\pi^{-}\Sigma^{+}$ mass spectrum at the energy around the $\bar{K}N$ threshold as shown in Fig. 6 of Ref. Xie:2018gbi . It should be pointed out that, the $\Sigma^{*}(1/2^{-})$ with a mass around 1430 MeV could be dynamically generated from the meson-baryon interaction, and appears as a weakly bound state or a virtual state, which is reflected as a strong cusp structure around the $\bar{K}N$ threshold Roca:2013cca . Meanwhile, the branching ratio for the $\Sigma^{*}(1/2^{-})$ signal is predicted to be of the order of $10^{-4}$ Xie:2018gbi . In the future, Belle II, BESIII, and the future Super Tau-Charm Factory (STCF) experiments could test such predictions.
@@ Acknowledgements
We would like to thank Eulogio Oset for useful comments. LSG and JJX are partly supported by the National Key R&D Program of China under Grant No. 2023YFA1606700. EW is partly supported by the National Key R&D Program of China under Grant No. 2024YFE0105200. This work is supported by the Natural Science Foundation of Henan under Grant No. 232300421140 and No. 222300420554, the National Natural Science Foundation of China under Grant Nos. 12475086, 12192263, 12205075, 12175239, 12221005, 12075288, 12361141819, and also by the Open Project of Guangxi Key Laboratory of Nuclear Physics and Nuclear Technology, No. NLK2021-08, and also by the Central Government Guidance Funds for Local Scientific and Technological Development, China (No. Guike ZY22096024), and also by the National Key Research and Development Program of China under Contracts 2020YFA0406400, and also by the Chinese Academy of Sciences under Grant No. YSBR-101, and also by the Youth Innovation Promotion Association CAS.